OpenStudy (anonymous):

$1000 was invested altogether. The first investment yielded a 15% profit, but the second investment yielded a 10% loss. Altogether this was equivalent to a 6% loss. How Much was in each of these investments?

OpenStudy (anonymous):

x + y = 1000 (1.15)x + (0.90)y = 940 Two equations in two unknowns. Just solve for x and y now. First equation is self-explanatory. In the second equation, (1.15)x is the expression for a 15% profit on the first investment. Similar for second term. 940 shows a 6% loss on 1000. Easy to solve at this point now that the equations are set up.

OpenStudy (anonymous):

Thanks. Can I ask you another question too? I need to know if I set it up correctly.

OpenStudy (anonymous):

Would it be, 1.15x for the first invest or .15x ?

OpenStudy (anonymous):

ok, but did you see what I did for the question above? No problems understanding?

OpenStudy (anonymous):

Kind of...

OpenStudy (anonymous):

where do you get the .9 from though

OpenStudy (anonymous):

I get everything except where you get the .9y, and why it is 1.15x as opposed to .15x

OpenStudy (anonymous):

You'll have to get that methodology down cold before you are able to do these on your own. 0.90y is the expression for a 10% loss because 90% of the investment remains.

OpenStudy (anonymous):

Whew! Back now, (for now!)

OpenStudy (anonymous):

Ok I understand now actually. It's common sense just not seeing it.

OpenStudy (anonymous):

1.15 because you retain the first investment (1) and show a profit (0.15), so principal + profit gives the 1.15

OpenStudy (anonymous):

These questions can be hard at first, but they will get easier with practice.

OpenStudy (anonymous):

Ahh I see. Are you going to be online?

OpenStudy (anonymous):

I'm going to see help with 3 more.

OpenStudy (anonymous):

Just post new question in a new thread. I'll look for you.

OpenStudy (anonymous):

I just need to know if I set up the equation right, should I make a thread for that?

OpenStudy (anonymous):

I can look here if you like.

OpenStudy (anonymous):

Ok. I'll make it fast. Thanks.

OpenStudy (anonymous):

CLyde has 900 invest at 8%. How much money must Clyde invest at 3% so that he will assume a 6% return altogether on the two investments.

OpenStudy (anonymous):

I put this equation...

OpenStudy (anonymous):

.06 = 900(.08) + x (.03)

OpenStudy (anonymous):

I came up with: 972 + (1.03)x = (900 + x)(1.06). That will work. I still have to check your method out. But first, I'm going to solve mine for x.

OpenStudy (anonymous):

Did you get 972 by multiplying 900 by .08

OpenStudy (anonymous):

Not exactly. I muliplied it by 1.08. Principal + profit

OpenStudy (anonymous):

By the way, my equation worked because I took my answer and went through the problem statement. You should get 600 for the principal to be invested at .03

OpenStudy (anonymous):

I believe you. I just need help setting it up.

OpenStudy (anonymous):

why do you add 1 to the left of the decimals. i'm confused on what you're referring to as the principal.

OpenStudy (anonymous):

We can either stick with my method and show why it works, and more importantly, get you into being able to set it up like that, or we can look at your equation and see what went wrong. I sort of suggest just looking at what I did and we can go through it more closely.

OpenStudy (anonymous):

Principal is what I'm calling the amount to be invested.

OpenStudy (anonymous):

lets do that.

OpenStudy (anonymous):

I add 1 to the decimal to show the "afterwards" on the investment, when there is a profit. When there is a loss, I subtract the loss percentage.

OpenStudy (anonymous):

ultimately saving time... good idea.

OpenStudy (anonymous):

So, starting with this last of the two problems, he invests 900 at 8%, so we know that from that 900, he will end up with 972 because he retains the 900 and gets a profit 900 x (1.08) = 972. 72 is the profit from 900 x 0.08. Okay so far?

OpenStudy (anonymous):

so why is it (at the end of the equation) multiplied by 1.06 as opposed to added.

OpenStudy (anonymous):

oh nevermind. because its the 6% of the return

OpenStudy (anonymous):

im with u

OpenStudy (anonymous):

i'll let u continue until the answer, then ask questions

OpenStudy (anonymous):

Well, we'll get to that, but that's not the next step. Gotta do this in order or it won't make sense. So, you're ok on the 972?

OpenStudy (anonymous):

yea. i under stand it.

OpenStudy (anonymous):

the 1 on the left represents the original investment.

OpenStudy (anonymous):

(left of decimal) yea I get that part. thank u alot.

OpenStudy (anonymous):

same rule applies for the rest because they're all an increase of return? right?

OpenStudy (anonymous):

Now, about the other investment, we don't know what it is yet, so we'll just call it x, but we do know that we will get a 3% profit from it. We will also get our investment back, so we will end up from the x with (1 + 0.03)x at the end of the investment period. Same as saying (1.03)x. Yes, same rule, just saw your last question.

OpenStudy (anonymous):

so the reason you put the 900 on the right side of the = is because it represents the original investment?

OpenStudy (anonymous):

So, he ends up with 900(1.08) + (1.03)x when all is said and done and the investment period is over. That is simplified to 972 + (1.03)x. Answering your last question, 972 is on the same side as the (1.03)x and represents the total amount he will have.

OpenStudy (anonymous):

That amount is the "afterward". What he starts with is 900 + x, two separate investments that he will look at together as far as Principal or amount to be invested as a "package" (conceptually). He wants that 900 + x as a hypothetical (for comparison purposes) amount invested at 6% to equal the "afterward" . So, he sets the hypothetical (900 + x)(1.06) as equal to the 972 + (1.03)x. Now we have one equation in one variable which is easier to solve than the first problem.

OpenStudy (anonymous):

So, that's basically it. You can feel free to ask questions about it, but that's basically all that I did.

OpenStudy (anonymous):

Okay, I see how you did that. Now, with any investment problems, you can use this same method. You just have to do the math for the difference in the investment, and represent properly if its a loss or gain... right?

OpenStudy (anonymous):

Bear in mind that though these two questions were investment problems, the way we eventually solved each was very different. There were some similarities in thinking about "before and after" and that profit gave multiplying principal by 1 + % returned (and loss by 1 - % lost). But notice how the first problem required 2 equations in 2 unknowns, and the second problem required 1 equation in 1 unknown.

OpenStudy (anonymous):

Right, that is just with what their asking. But essentially for investments, and when they say gain or loss you can use the 1 + .05 or whatever the percent is. or 1 - .05 if it's a loss.

OpenStudy (anonymous):

i have another question. where i need to see if I set it up correctly. can I ask you here? i dont want someone to answer with a different method and confuse me. if not its ok.

OpenStudy (anonymous):

Yes, that will be a great help conceptually and mathematically. And a little bit of a mental shortcut.

OpenStudy (anonymous):

On last question, sure, go ahead. But before we get into it, I just want to say that, once we are all done, you might want to go back and review this whole thread at some time, because I think it will help.

OpenStudy (anonymous):

Great. Got that embedded now, excellent method.

OpenStudy (anonymous):

Yea. I'm new to this site, do the chat logs stay saved?

OpenStudy (anonymous):

You click on your name near the top right and that will open a second window. Then you go to where you have "questions asked". It will be there but I don't know how long it is saved.

OpenStudy (anonymous):

There is a total of $7000 in two investments. The annual income from the 11% investment is $50 more than the annual income from the 9% investment. How much is invested at each rate? I did two equations. x + y = 7000 .11x = .9y + 50 correct?

OpenStudy (anonymous):

I 've been here only under 2 weeks, so maybe they go away after a while.

OpenStudy (anonymous):

Oh okay. Thank you for your help. It's really appreciated.

OpenStudy (anonymous):

On that second equation, be careful. It would be .09, not .9 .9 is 90%

OpenStudy (anonymous):

Ahhhhh... I need to observe! How was the actual set up, minus that?

OpenStudy (anonymous):

On a first look, it looks good, but I always work it through before giving a final answer. So, I'm going through it now.

OpenStudy (anonymous):

Ok. great. I primarily need to get the steps of setting up the equation more, as opposed to doing the actual arithmetic behind it. The setup is very difficult to me.

OpenStudy (anonymous):

There's something wrong here. Is there something left out of the problem statement? There is reference to annual income. Is there a investment length period overall?

OpenStudy (anonymous):

As far as the question, that is all that is provided.

OpenStudy (anonymous):

OK, then it works out just fine. After doing the math, you should get 3600 for the .09 investment and 3400 for the .11 investment.

OpenStudy (anonymous):

So, good job! You came up with those equations on your own!

OpenStudy (anonymous):

Yup. It takes a lot of re-practicing for these. Not as easy as it was in high school lol. May I ask you another? And is there any more feedback I can give you to boost your profile on this site? Nothing but A+++ feedback lol. (If you're familiar with eBay)

OpenStudy (anonymous):

I don't go in so much for the recognition. I joined because I had one stumper of a question which I eventually figured out myself. Then I hung around because I like helping out and I like math. So, you don't have to give any kudos, really. There is some provision somewhere for what they call fan testimonial, but I don't know too much about it. You can ask another question, but then I have to go.

OpenStudy (anonymous):

Alright I'll look into that. Thank you again, I'll make this last one fast...

OpenStudy (anonymous):

$5000 is in three investments. Twice as much is invested at 8% as at 6% and the remainder is invest at 9%. If the annual return is $390, how much is invested at each rate?

OpenStudy (anonymous):

You can take your time, don't rush, these things can get tricky if rushed.

OpenStudy (anonymous):

I understand, but I don't want to hinder your schedule. Now, I'm really confused on the "remainder" into 9% I don't know how to represent that. So far i've came up with x + y + z = 5000 and (.08)x and y(.06)(2)

OpenStudy (anonymous):

So you can get an idea of what I understand, I'll let you explain and I'll ask questions after you're done.

OpenStudy (anonymous):

Ok, I've got it.

OpenStudy (anonymous):

Mathematically, it's 3 equations in 3 unknowns which on the surface makes this problem at first appear harder than it really is. It really simplifies down to 2 equations in 2 unknowns which are solved first, and then the last equation in 3 unknowns is actually very easily solved. Ok, so much for the math analysis, we'll get to it now.

OpenStudy (anonymous):

You're off to a good start with x + y + z = 5000, so let's go with those variables and that as the first equation. Also, let's call x the amount we invest at .06, y the amount at .08, and z the amount at .09.

OpenStudy (anonymous):

Im with you...

OpenStudy (anonymous):

"Twice as much is invested at 8% as at 6%" is a big key to simplifying. It tells us that y = 2x. So, equation 2 becomes x + 2x + z = 5000 and simplifies to 3x + z = 5000. That's going to be a big help.

OpenStudy (anonymous):

It helps simplify equation 3 tremendously. Equation 3 starts out as (0.06)x + (0.08)y + (0.09)z = 390 and ends up as (0.06)x + (0.08)(2x) + (0.09)z = 390 which ends up as (0.22)x + (0.09)z = 390 . So, this eventual equation 3 ends up as being in the same two variables as equation 2. So, you solve these two equations first for x and z, and then simply plug those into equation 1.

OpenStudy (anonymous):

So, all that's left is to do the math, but that's easy now. If you want, I can try to get the actual numbers.

OpenStudy (anonymous):

now to plugin the 3 equation to the 1st. I would have to solve for x or z. in the second equation i solved for Z, so I would have to solve for X. i this case i will subtract the .09z from both sides, then divide the .22 from both sides to isolate the X ?

OpenStudy (anonymous):

You should get x = 1200, y = 2400, and z = 1400.

OpenStudy (anonymous):

The method is to solve equations 2 and 3 together for x and z. You will get actual numbers for x and z. Take the x and z (not equations 2 or 3) and put the numerical x and z into equation 1 and you will get y. Or just remember that y is twice x.

OpenStudy (anonymous):

so .22x + y + .09z = 5000 ?

OpenStudy (anonymous):

Remember, equation 2 is 3x + z = 5000. Equation 3 is (0.22)x + (0.09)z = 390. Easy way to solve that is to take equation 2 and that z = 5000 - 3x. Take THAT z and put it into the z in equation 3. No "y" in your equation above.

OpenStudy (anonymous):

so. (0.22)x + (0.09)(5000 - 3x) = 390 ?

OpenStudy (anonymous):

Equation 3 will become (0.22)x + (0.09)(5000 - 3x) = 390. NOW, you've got one equation in one variable, so you can solve for x and you should get 1200 for x. Just saw your post. Looks like you figured out that part that I just mentioned. You're getting it.

OpenStudy (anonymous):

Yup. Thanks to you. I'll need to review some of these at my lab tomorrow.

OpenStudy (anonymous):

They're actually fun once you get the hang of it. After you review, you'll be surprised at how easy it gets.

OpenStudy (anonymous):

Yea I can see how it becomes easier. It's all the set up, the arithmetic is relatively easy

OpenStudy (anonymous):

Yes. Well, take care now. It's been great working with you. And good luck in all your studies. Bye for now.

OpenStudy (anonymous):

Thank you for you're time. Have a great evening...